### v2

@@ A.3 Lexicon-based approach to measure sentiment from news articles
Another approach to measure sentiment regarding economic outlook conveyed by news articles is the lexicon-based approach where one relies on a predefined list of words that are associated with positive, negative, or neutral sentiment (see, e.g., loughran:2011). Occurrences of these words are counted in a text and a sentiment score based on these occurrences is assigned. We use the lexicon developed by barbaglia:2025. The lexicon is in English and consists of more than six thousand words. We translated and adapted it to German and applied it to the German news articles. The results of the lexicon-based approach are not very promising. Simple word-count techniques face challenges in recognizing context and detecting negations in texts, for example “not happy” (see, e.g., binsbergen:2024). We could translate the lexicon to French as well, but we refrain from the additional effort as the German articles are a clear majority in our sample and in an additional computing exercise, we observed that NEOS with and without the French articles looks very similar. Thus, the French articles do not seem to contain much additional information for sentiment, and this should also hold applying other approaches such as the lexicon-based one.
A.4 Forecast improvement during crises
Table 1 shows that including NEOS in the regression model in Equation (2) improves the accuracy of GDP growth forecasts compared to the AR(1) model. To reveal whether the improvements are due to consistent gains across many quarters or due to large gains during a few key episodes, for the case of forecasting with horizon $h=0$, we first compute the squared errors for each quarter made by both the regression model in Equation (2) and the AR(1) model using the same pseudo-out-of-sample expanding window framework as for Table 1. Second, we subtract the squared errors of the AR(1) model from those of the regression model (2) and plot the cumulative sum of these differences in Figure 13. Declines in the cumulative series indicate periods in which the regression model (2) outperforms the AR(1) model.
Figure 13: Cumulative forecast improvements during key economic crises.

### v3

@@ A.3 EPU and lexicon-based approach to measure sentiment from news articles
b) To construct an alternative text-based sentiment indicator regarding the economic outlook we apply a lexicon-based approach where one relies on a predefined list of words that are associated with positive, negative, or neutral sentiment (see, e.g., Loughran and Mcdonald (2011)). Occurrences of these words are counted in a text and a sentiment score based on these occurrences is assigned. We use the lexicon developed by Barbaglia et al. (2025). The lexicon is in English and consists of more than six thousand words. We translated and adapted it to German and applied it to the German news articles. [^9] All articles in German that enter the computation of NEOS are used. The results of the lexicon-based approach are not very promising. Simple word-count techniques face challenges in recognizing context and detecting negations in texts, for example, “not happy” (see, e.g., van Binsbergen et al. (2024)). We could also translate the lexicon to French, but we refrain from the additional effort as the German articles are a clear majority in our sample and in an additional computing exercise, we observed that NEOS with and without the French articles looks very similar. Thus, the French articles do not seem to contain much additional information for sentiment, and this should also hold applying other approaches such as the lexicon-based one.
A.4 Forecast evaluation for subperiods
Table A6: Forecasting Swiss year-on-year GDP growth from Q1 2009 to Q1 2025
Indicator $h=0$ $h=1$ $h=2$
NEOS 0.94 (0.18) 1.00 (0.51) 0.93 (0.10)
NEOS (first 7 days in the third month) 0.93 (0.17) 0.99 (0.44) 0.93 (0.09)
NEOS (first 14 days in the third month) 0.94 (0.21) 0.98 (0.43) 0.94 (0.12)
NEOS (first 21 days in the third month) 0.95 (0.25) 0.98 (0.41) 0.94 (0.13)
Lexicon-based approach 1.20 (0.95) 1.11 (0.74) 1.05 (0.68)
EPU for Switzerland 1.12 (0.86) 1.11 (0.89) 1.01 (0.52)
Manufacturing PMI for Switzerland 1.03 (0.61) 1.01 (0.53) 0.90 (0.20)
KOF Business Situation Indicator 1.06 (0.86) 1.00 (0.67) 1.00 (0.48)
SECO Consumer Sentiment Index 1.20 (0.79) 1.07 (0.76) 1.00 (0.52)
MAE ratios with the Diebold–Mariano p-values in parentheses. The best MAE values are in bold.
Table A7: Forecasting Swiss year-on-year GDP growth from Q1 2014 to Q1 2025
Indicator $h=0$ $h=1$ $h=2$
NEOS 0.94 (0.20) 0.91 (0.09) 0.89 (0.04)
NEOS (first 7 days in the third month) 0.94 (0.22) 0.91 (0.11) 0.88 (0.04)
NEOS (first 14 days in the third month) 0.95 (0.28) 0.91 (0.13) 0.88 (0.06)
NEOS (first 21 days in the third month) 0.97 (0.33) 0.90 (0.12) 0.87 (0.06)
Lexicon-based approach 1.29 (0.97) 1.06 (0.60) 1.05 (0.62)
EPU for Switzerland 0.99 (0.49) 1.04 (0.88) 0.99 (0.45)
Manufacturing PMI for Switzerland 1.05 (0.64) 0.84 (0.24) 0.87 (0.29)
KOF Business Situation Indicator 1.05 (0.63) 1.01 (0.63) 0.97 (0.30)
SECO Consumer Sentiment Index 1.10 (0.62) 1.00 (0.51) 1.05 (0.70)
Service PMI for Switzerland 0.86 (0.12) 0.94 (0.16) 0.93 (0.09)
MAE ratios with the Diebold–Mariano p-values in parentheses. The best MAE values are in bold.
A.5 Cumulative absolute error differences for all horizons
Figure 14: Cumulative forecast improvements compared to an AR(1) model for forecast horizons $h=0$ (top), $h=1$ (middle) and $h=2$ (bottom).
A.6 Accuracy vs. precision for different candidate models to classify relevant articles
Figure 15: Accuracy vs. precision for different candidate models to classify relevant articles. We use the abbreviations random forest (RF), L2-regularized logistic regression (LR), and ReLU feed-forward neural network (NN). (RF + NN)/2 is the equally-weighted ensemble. Every point in the plot corresponds to a particular threshold $\tau\in[0.5,0.99]$. We assign label 1 whenever the predicted probability is larger than $\tau$.